OpenStudy (anonymous):

Find the circumference of the circle. Use p » 3.14

OpenStudy (anonymous):



OpenStudy (lgbasallote):

\[C = 2 \times 3.14 \times r\] does that help?

OpenStudy (anonymous):

no sorry i dont understand

OpenStudy (lgbasallote):

r means radius. in this case radius is 3 so circumference is just \[2 \times 3.14 \times 3\] now do you understand?

OpenStudy (anonymous):

yes thanks

OpenStudy (lgbasallote):

welcome